# Finite-time blowup in a five- and higher-dimensional chemotaxis model with indirect signal production

### v2

@@ 1. Introduction
The mass identity (2.4) follows directly from integrating the first and last equations in (1.3) over $\Omega$. Integrating the second equation yields
$\int_{\Omega}v(x,t)=e^{-t}\int_{\Omega}v_{0}+(1-e^{-t})\int_{\Omega}u_{0}\leq\max\left\{\int_{\Omega}v_{0},\int_{\Omega}u_{0}\right\}\quad\text{for all }t\in(0,T_{\max}).$
By applying well-known $W^{1,p}$ regularity theory [1] to the elliptic boundary value problem $-\Delta w+w=u$, we obtain that for each $q\in[1,n/(n-1))$, there exists a constant $C_{q}>0$ such that
$\|w\|_{W^{1,q}(\Omega)}\leq C_{q}\|u\|_{L^{1}(\Omega)}$
holds for all $t\in(0,T_{\max})$. This implies (2.6) in view of (2.4). Thus, $w\in L^{p}(\Omega)$ for $p\in[1,n/(n-2))$ by the Sobolev embedding theorem. Moreover, for any $p\in[1,n/(n-2))$ one can select $q\in[1,n/(n-1))$ such that
$-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)>-1.$ (2.8)
[…]
$\|\nabla v_{x_{i}}\|_{L^{p}(\Omega)}$ $\leq C\|v_{0x_{i}}\|_{W^{1,p}(\Omega)}$
$\quad+C\int_{0}^{t}\left(1+(t-s)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)}\right)e^{-(t-s)}\;\mathrm{d}s\sup_{s\in(0,t)}\|w_{x_{i}}(\cdot,s)\|_{L^{q}(\Omega)}$
$\leq C\|v_{0}\|_{W^{2,2}(\Omega)}+CC_{q}m+CC_{q}m\Gamma\left(\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)\right),\quad i\in\{1,2,\cdots,n\},$
holds for all $t\in(0,T_{\max})$, $p\in[1,n/(n-2))$, where the Euler’s Gamma function $\Gamma$ is well-defined due to the choices of parameters in (2.8). According to (2.5), this implies that $v\in W^{2,p}(\Omega)$ for all $t\in(0,T_{\max})$, as well as (2.7), since $v_{t}=\Delta v-v+w\in L^{p}(\Omega)$ for all $t\in(0,T_{\max})$. ∎
3. Pointwise estimates for chemical concentration
This section is devoted to the deduction of uniform-in-time pointwise estimates for $v$ and $w$, based on Lemma 2.3 with additional symmetry assumptions. For convenience, we shall switch to the radial notation without any further comment, e.g., writing $v(r,t)$ instead of $v(x,t)$ when appropriate. Throughout the sequel, we assume that the radius $R$ of the ball $\Omega$ is fixed. Consequently, it will not be explicitly stated that the generic constants $C(p),C(m,M,B,\kappa),\ldots$ may vary from line to line, and may additionally depend on $R$.
@@ 3. Pointwise estimates for chemical concentration
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ with some $R>0$ and $n\geq 4$. If $(u,v,w)$ is a classical solution of (1.1) associated with radially symmetric initial datum $(u_{0},v_{0})$ as given in Proposition 2.1, then there exists a constant $C>0$ such that
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq Cm$ (3.1)
for all $t\in(0,T_{\max})$; and then for each $\beta>n-2$, there exists $C(\beta)>0$ such that
$\||x|^{\beta-1}v\|_{W^{1,\infty}(\Omega)}\leq C(\beta)(m+\|v_{0}\|_{W^{2,2}(\Omega)})$ (3.2)
for all $t\in(0,T_{\max})$.
@@
The weighted uniform estimate (3.1) follows directly from Lemma 3.2. The estimates in Lemma 3.3 yield (3.2). ∎
The mass identity (2.4) permits us to gain pointwise information about $w$.
@@
Let $w$ be as defined in Proposition 3.1. Then there exists a constant $C>0$ such that the nonnegative function $w$ satisfies
$w+|x\cdot\nabla w|\leq\frac{Cm}{|x|^{n-2}},\quad|x|\in(0,R],$ (3.3)
for all $t\in(0,T_{\max})$.
@@
Integrating the last equation of (1.1), $ur^{n-1}=wr^{n-1}-(w_{r}r^{n-1})_{r}$ over the interval $(0,\rho)\subset(0,R)$ yields
@@
for all $t\in(0,T_{\max})$, where the mass identity (2.4) has been used. Here and henceforth, $\omega_{n}$ denotes the surface area of the unit ball in $\mathbb{R}^{n}$. Choose $r_{0}(t)\in(R/2,R)$ such that
@@
$w=w(r_{0})+\int_{r_{0}}^{r}w_{r}(\eta)\;\mathrm{d}\eta\leq\frac{m}{|B_{R}\setminus B_{R/2}|}+\frac{2\omega_{n}^{-1}m}{n-2}\left(r^{2-n}+r_{0}^{2-n}\right),$
for all $t\in(0,T_{\max})$. Setting a suitably large constant $C(m)>0$, such as
$C(m):=\frac{m}{|B_{R}\setminus B_{R/2}|}+\frac{2^{n}m}{\omega_{n}},$
@@
Following the standard approach in [47], we can deduce pointwise estimates for the gradient of $v$ from (2.7).
@@
Let $v$ be as defined in Proposition 3.1. Then for any fixed $\beta>n-2$,
@@
holds for some $C(\beta)>0$.
@@
According to (2.7), we know that $v\in W^{2,p}(\Omega)$ for $p\in[1,n/(n-2))$. Hölder inequality entails
@@
for all $r\in(0,R)$, $p\in(1,n/(n-2))$ and $t\in(0,T_{\max})$, where $r_{1}(t)\in(R/2,R)$ is chosen such that
@@
$\leq\frac{\|\nabla v\|_{L^{1}(\Omega)}}{|B_{R}\setminus B_{R/2}|}+C_{s}\|v\|_{W^{2,p}(\Omega)}\left(r^{-\frac{n-p}{p}}+r_{1}^{-\frac{n-p}{p}}\right),$
for all $r\in(0,R)$, $p\in(1,n/(n-2))$ and $t\in(0,T_{\max})$, with some $C_{s}>0$ inferred from the Sobolev inequality. Since $\beta>n-2$, we can put
$p=\max\left\{\frac{n}{n-1},\frac{n}{\beta}\right\}<\frac{n}{n-2}$
and by applying an analogous argument to the one used in Lemma 3.2, we conclude the proof. ∎
4. Energy bounded below by sublinear dissipation rate
The goal of this section is to establish a lower bound for the energy functional $\mathcal{F}$ in terms of a power of the dissipation rate $\mathcal{D}$, as discussed in the introduction, where $\mathcal{F}$ and $\mathcal{D}$ are defined in (2.3).
To relate the energy $\mathcal{F}$ to the dissipation rate $\mathcal{D}$, we introduce the following quantities
$f:=-\Delta v+v-w$ (4.1)
@@ 4. Energy bounded below by sublinear dissipation rate
$g:=\left(\frac{\nabla u}{\sqrt{u}}-\sqrt{u}\nabla v\right)\cdot\frac{x}{|x|},\quad x\in\overline{\Omega}\setminus\{0\},$ (4.2)
where $w=(-\Delta_{N}+1)^{-1}u$ is the unique classical solution to the Neumann elliptic problem (1.4). Then the energy functional $\mathcal{F}$ and the dissipation rate $\mathcal{D}$ are given by
$\mathcal{F}(u,v)=\int_{\Omega}(u\log u-uv)+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v|^{2}\quad\text{and}$ (4.3)
$\mathcal{D}(u,v)=\int_{\Omega}\left(\left|\nabla f\right|^{2}+\left|f\right|^{2}\right)+\int_{\Omega}g^{2}.$
@@
$\int_{\Omega}u=m,\quad\int_{\Omega}v\leq M\quad\text{and }L^{1}\text{-bound}\quad\int_{\Omega}|f|\leq B,$ (4.4)
@@
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq M\quad\text{and}\quad\||x|^{\kappa-1}v\|_{W^{1,\infty}(\Omega)}\leq B,$ (4.5)
@@
where $\theta\in(0,1)$ and $C(m,M,B,\kappa)>0$. Here $\mathcal{S}(m,M,B,\kappa)$ is the set of functions defined as
@@
[…]
We begin with the following straightforward observation. It can be seen that controlling the energy functional $\mathcal{F}$ from below is equivalent to bounding the mixed term $\int_{\Omega}uv$ from above.
@@
For all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$,
$-\mathcal{F}(u,v)\leq\int_{\Omega}uv+\frac{\omega_{n}R^{n}}{e}$ (4.6)
holds, where $\omega_{n}$ is the surface area of the unit ball in $\mathbb{R}^{n}$ and $R$ denotes the radius of the ball $\Omega$.
@@
This result follows directly from the basic inequality $s\ln s\geq-e^{-1}$ valid for all $s>0$. ∎
Motivated by Winkler [47], we shall estimate the integral $\int_{\Omega}uv$ by separately considering the inner region $B_{\rho}$ and the outer region $\Omega\setminus B_{\rho}$. Additionally, we incorporate decoupling techniques from Cao and Fuest [4], which assist us to convert the mixed term into an interior $L^{2}$–estimate for the Laplacian of the chemical concentration $v$.
@@
There exists a constant $C(\kappa)>0$ such that
$\int_{\Omega}uv\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}+C(\kappa)(m+M)B\rho^{2-\kappa},$
@@
For convenience, we introduce a cut-off function defined as
$\phi^{(\rho)}(x):=\phi(x/\rho),\quad x\in\mathbb{R}^{n},\quad\rho\in(0,R),$
where $\phi(x)\in C_{0}^{\infty}(\mathbb{R}^{n})$ is a nonnegative and radially symmetric function such that
$\phi(x)=\begin{cases}1,&\text{if }x\in B_{1/2},\\ 0,&\text{if }x\in\mathbb{R}^{n}\setminus B_{1}.\end{cases}$
It follows that
$\rho\|\nabla\phi^{(\rho)}\|_{L^{\infty}(\mathbb{R}^{n})}+\rho^{2}\|\Delta\phi^{(\rho)}\|_{L^{\infty}(\mathbb{R}^{n})}\leq C_{\phi}:=\|\nabla\phi\|_{L^{\infty}(\mathbb{R}^{n})}+\|\Delta\phi\|_{L^{\infty}(\mathbb{R}^{n})}$
for each $\rho\in(0,R)$. Testing $u=-\Delta w+w$ by $v\phi^{(\rho)}$, we compute upon integration by parts,
$\int_{B_{\rho/2}}uv$ $\leq\int_{\Omega}uv\phi^{(\rho)}=-\int_{\Omega}\Delta wv\phi^{(\rho)}+\int_{\Omega}wv\phi^{(\rho)}=\int_{\Omega}w(-\Delta(v\phi^{(\rho)})+v\phi^{(\rho)})$ (4.7)
$=\int_{\Omega}(-\Delta v+v)w\phi^{(\rho)}-\int_{\Omega}w(2\nabla v\cdot\nabla\phi^{(\rho)}+v\Delta\phi^{(\rho)})=:J_{1}-J_{2}.$
In view of $w=-\Delta v+v-f$ specified in (4.1), we estimate by Young inequality
$J_{1}$ $=\int_{\Omega}(-\Delta v+v)^{2}\phi^{(\rho)}+\int_{\Omega}(\Delta v-v)f\phi^{(\rho)}\leq\frac{3}{2}\|\Delta v-v\|_{L^{2}(B_{\rho})}^{2}+\frac{1}{2}\|f\|_{L^{2}(B_{\rho})}^{2}$ (4.8)
$\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}.$
Recalling pointwise information from (4.5), we estimate
$-J_{2}$ $\leq 2MBC_{\phi}\int_{B_{\rho}\setminus B_{\rho/2}}|x|^{2-n}(|x|^{1-\kappa}|x|^{-1}+|x|^{2-\kappa}|x|^{-2})$ (4.9)
$=2MBC_{\phi}\omega_{n}\int_{\rho/2}^{\rho}r^{1-\kappa}\;\mathrm{d}r\leq 2^{\kappa-1}\omega_{n}MBC_{\phi}\rho^{2-\kappa}.$
Using the pointwise information from (4.5), we also estimate the integral in the annular region
$\int_{\Omega\backslash B_{\rho/2}}uv\leq\sup_{r\in\left(\rho/2,R\right)}v(r)\cdot\int_{\Omega\setminus B_{\rho/2}}u\leq B\sup_{r\in\left(\rho/2,R\right)}r^{2-\kappa}\cdot\int_{\Omega}u=2^{\kappa-2}Bm\rho^{2-\kappa}$ (4.10)
holds for all $(u,v)\in\mathcal{S}$ and $\rho\in(0,R)$. Summing (4.7)–(4.10) yields the desired estimate
$\int_{\Omega}uv\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}+(2^{\kappa-1}\omega_{n}MC_{\phi}+2^{\kappa-2}m)B\rho^{2-\kappa},$
which completes the proof. ∎
To utilize the information from (4.2), we adapt the Pohožaev-type test function $|x|^{n-2}x\cdot\nabla v$, originally proposed by Winkler [45, 46, 47], in order to bound the second-order integral from above by terms from dissipation rate and zeroth-order signal terms. We believe this test scheme has been widely employed in the analysis of elliptic problems [2].
@@
Let $n\geq 5$. Then there exists a constant $C_{0}>0$ such that for each $\rho\in(0,R)$,
$\frac{1}{8}\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+\frac{3}{4}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}$ $\leq 12\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2}+C_{0}\|f\|_{L^{2}(B_{\rho})}^{2}+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}$ (4.11)
$\quad+\|v\|_{L^{2}(B_{\rho})}^{2}+C_{0}B^{2}\rho^{n+2-2\kappa}+C_{0}BM\rho^{2-\kappa}+2m$
holds for all $(u,v)\in\mathcal{S}$.
@@
$-\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r+\int_{0}^{\eta}wv_{r}r^{\alpha}\;\mathrm{d}r$ $=\int_{0}^{\eta}(w_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r.$ (4.12)
@@
$\int_{0}^{\eta}(w_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r$ $=w_{r}(\eta)v_{r}(\eta)\eta^{\alpha}-\int_{0}^{\eta}w_{r}r^{n-1}(v_{r}r^{n-1})_{r}\;\mathrm{d}r$ (4.13)
$=w_{r}(\eta)v_{r}(\eta)\eta^{\alpha}-\int_{0}^{\eta}w_{r}yr^{\alpha}\;\mathrm{d}r,\quad\eta\in(0,R).$

### v3

@@ 2 Preliminaries
The mass identity (12) follows directly from integrating the first and last equations in (3) over $\Omega$. Integrating the second equation yields
$\int_{\Omega}v(x,t)=e^{-t}\int_{\Omega}v_{0}+(1-e^{-t})\int_{\Omega}u_{0}\leq\max\left\{\int_{\Omega}v_{0},\int_{\Omega}u_{0}\right\}\quad\text{for all }t\in(0,T_{\max}),$
which shows (13). By applying well-known $W^{1,p}$ regularity theory [33] to the elliptic boundary value problem $\Delta w+w=u$, we obtain that for each $q\in\Big[1,\frac{n}{n-1}\Big)$, there exists a constant $C(q)>0$ such that $\|w\|_{W^{1,q}(\Omega)}\leq C(q)\|u\|_{L^{1}(\Omega)}$ holds for all $t\in(0,T_{\max})$. This implies (14) by virtue of (12). Thus, $w\in L^{p}(\Omega)$ for $p\in\Big[1,\frac{n}{n-2}\Big)$ by Sobolev embedding $W^{1,q}\hookrightarrow L^{\frac{nq}{n-q}}$ for $q\in\Big[1,\frac{n}{n-1}\Big)$. Moreover, for any $p\in\Big[1,\frac{n}{n-2}\Big)$ one can select $q\in\Big[1,\frac{n}{n-1}\Big)$ such that
$\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)>-1.$ (16)
$L^{p}$$L^{q}$ estimates of the heat semigroup $e^{t\Delta}$ with homogeneous Dirichlet boundary conditions (cf. [34, Lemma 2.1] and [35, Lemma 2.4]) applied to $v_{x_{i}t}=\Delta v_{x_{i}}-v_{x_{i}}+w_{x_{i}}$ subject to $v_{x_{i}}=v_{r}x_{i}/r\equiv 0$ on $\partial\Omega$ for $i\in\{1,2,\ldots,n\}$, provide a constant $C_{d}>0$ such that
$\|\nabla v_{x_{i}}\|_{L^{p}(\Omega)}$ $\leq\|\nabla e^{t(\Delta-1)}v_{0x_{i}}\|_{L^{p}(\Omega)}+\int_{0}^{t}\|\nabla e^{(t-\tau)(\Delta-1)}w_{x_{i}}\|_{L^{p}(\Omega)}\;\mathrm{d}\tau$
$\leq|\Omega|^{1-p/2}\|\nabla e^{t(\Delta-1)}v_{0x_{i}}\|_{L^{2}(\Omega)}$
$\quad+C_{d}\sup_{\tau\in(0,t)}\|w_{x_{i}}(\cdot,\tau)\|_{L^{q}(\Omega)}\cdot\int_{0}^{t}\left(1+(t-\tau)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)}\right)e^{-(t-\tau)}\;\mathrm{d}\tau$
$\leq C_{d}\|v_{0}\|_{W^{2,2}(\Omega)}+C_{d}C(q)m\bigg(1+\int_{0}^{\infty}\tau^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)}e^{-\tau}\;\mathrm{d}\tau\bigg),$
holds for all $t\in(0,T_{\max})$, $p\in\Big[1,\frac{n}{n-2}\Big)$ and $i\in\{1,2,\ldots,n\}$, where the last integral is finite due to the choices of parameters in (16). According to (13), this implies that $v\in W^{2,p}(\Omega)$ for all $t\in(0,T_{\max})$, and (15) holds, since $v_{t}=\Delta v-v+w\in L^{p}(\Omega)$ for all $t\in(0,T_{\max})$. ∎
3 Pointwise estimates for chemical concentration
This section is devoted to the deduction of uniform-in-time pointwise estimates for $v$ and $w$, based on Lemma 2.3. Specifically, we show that symmetric functions from Sobolev spaces can be embedded into certain weighted Sobolev spaces, as established in [36]. For convenience, we shall switch to the radial notation without any further comment, e.g., writing $v(r,t)$ instead of $v(x,t)$ when appropriate. Throughout the sequel, we assume that the radius $R$ of the ball $\Omega$ is fixed. Consequently, it will not be explicitly stated that generic constants $C(p),C(m,M,B,\kappa),\ldots$ may vary from line to line, and additionally depend on $R$.
@@ 3 Pointwise estimates for chemical concentration
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ with some $R>0$ and $n\geq 4$. There exists a constant $C_{w}>0$ such that for any radially symmetric initial datum $(u_{0},v_{0})$ with (6) and (7), the corresponding classical solution $(u,v,w)$ of (1) satisfies
$\sup_{t\in(0,T_{\max})}\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq C_{w}m.$ (17)
For each $\beta>n-2$, there exists $C(\beta)>0$ such that
$\sup_{t\in(0,T_{\max})}\||x|^{\beta-1}v\|_{W^{1,\infty}(\Omega)}\leq C(\beta)(m+\|v_{0}\|_{W^{2,2}(\Omega)})$ (18)
holds whenever radially symmetric initial data $(u_{0},v_{0})$ comply with (6) and (7).
@@
The weighted uniform estimate (17) follows directly from Lemma 3.2. The estimates in Lemma 3.3 yield (18). ∎
The mass identity (12) permits us to gain pointwise information about $w$.
@@
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ with some $R>0$ and $n\geq 4$. Then there exists a constant $C_{w}>0$ such that
$w+|x\cdot\nabla w|\leq\frac{C_{w}m}{|x|^{n-2}},\quad|x|\in(0,R],\quad t\in(0,T_{\max})$ (19)
for all $w$ given in Proposition 3.1.
@@
Integrating the last equation in (1), $ur^{n-1}=wr^{n-1}-(w_{r}r^{n-1})_{r}$ over the interval $(0,r)\subset(0,R)$ yields
@@
for all $t\in(0,T_{\max})$, where the mass identity (12) has been used. Here and henceforth, $\omega_{n}$ denotes the surface area of the unit ball in $\mathbb{R}^{n}$. Choose $r_{0}(t)\in(R/2,R)$ such that
@@
$\begin{split}w=w(r_{0})+\int_{r_{0}}^{r}w_{r}(\eta)\;\mathrm{d}\eta&\leq\frac{m}{|B_{R}\setminus B_{R/2}|}+\frac{2\omega_{n}^{-1}m}{n-2}\left(r^{2-n}+r_{0}^{2-n}\right)\\ &\leq\frac{mR^{n-2}r^{2-n}}{|B_{R}\setminus B_{R/2}|}+\frac{2\omega_{n}^{-1}m}{n-2}\left(r^{2-n}+2^{n-2}r^{2-n}\right)\end{split}$
for all $t\in(0,T_{\max})$. Setting
$C_{w}:=\frac{R^{n-2}}{|B_{R}\setminus B_{R/2}|}+\frac{2^{n}}{\omega_{n}},$
@@
Following the standard approach, e.g. in [37, 6], we can deduce pointwise estimates for the gradient of $v$ from (15). See [36, Theorem 1.1] for generalized results concerning Sobolev imbedding in the spaces of radial functions.
@@
For any $\beta>n-2$, there exists $C(\beta)>0$ such that
@@
holds for all $v$ given in Proposition 3.1.
@@
Since $\beta>n-2$, we can put
$p:=\max\bigg\{\frac{n}{n-1},\frac{n}{\beta}\bigg\}\in\bigg(1,\frac{n}{n-2}\bigg).$ (20)
Hölder inequality entails
@@
for all $r\in(0,R)$ and $t\in(0,T_{\max})$, where $r_{1}(t)\in(R/2,R)$ is chosen such that
@@
$\leq C_{s}\|v\|_{W^{2,p}(\Omega)}\left(r^{-\frac{n-p}{p}}+r_{1}^{-\frac{n-p}{p}}+\frac{1}{|B_{R}\setminus B_{R/2}|}\right),$
for all $r\in(0,R)$ and $t\in(0,T_{\max})$, with some $C_{s}>0$ inferred from Sobolev imbedding $W^{2,p}(\Omega)\hookrightarrow W^{1,\frac{np}{n-p}}(\Omega)\hookrightarrow W^{1,1}(\Omega)$. According to (15) and (20), the desired pointwise estimates follow from an analogous argument used in Lemma 3.2. ∎
4 Energy bounded below by sublinear dissipation rate
The goal of this section is to establish a lower bound for the energy functional $\mathcal{F}$ in terms of a power of the dissipation rate $\mathcal{D}$, built upon the functional framework originally proposed by Winkler [6], as discussed in the introduction, where $\mathcal{F}$ and $\mathcal{D}$ are defined in (11).
To relate the energy $\mathcal{F}$ to the dissipation rate $\mathcal{D}$ under symmetry assumptions, we introduce the following quantities
$f:=-\Delta v+v-w$ (21)
@@ 4 Energy bounded below by sublinear dissipation rate
$g:=\left(\frac{\nabla u}{\sqrt{u}}-\sqrt{u}\nabla v\right)\cdot\frac{x}{|x|},\quad x\in\overline{\Omega}\setminus\{0\},$ (22)
where
$w:=(-\Delta_{N}+1)^{-1}u$ (23)
is the unique classical solution to the Neumann elliptic problem (4). Then the energy functional $\mathcal{F}$ and the dissipation rate $\mathcal{D}$ are given by
$\begin{split}&\mathcal{F}(u,v)=\int_{\Omega}(u\log u-uv)+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v|^{2}\quad\text{and}\\ &\mathcal{D}(u,v)=\int_{\Omega}\left(\left|\nabla f\right|^{2}+\left|f\right|^{2}\right)+\int_{\Omega}g^{2}.\end{split}$ (24)
@@
$\int_{\Omega}u=m,\quad\int_{\Omega}v\leq M\quad\text{and }L^{1}\text{-bound}\quad\int_{\Omega}|f|\leq B,$ (25)
@@
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq M\quad\text{and}\quad\||x|^{\kappa-1}v\|_{W^{1,\infty}(\Omega)}\leq B,$ (26)
@@
where $\theta\in(0,1)$ and $C(m,M,B,\kappa)>0$. Here and below, $\mathcal{S}(m,M,B,\kappa)$ is the set of function pairs defined as
$\mathcal{S}(m,M,B,\kappa)$ $=\{(u,v)\in C^{1}(\bar{\Omega})\times C^{3}(\bar{\Omega})\mid u\text{ and }v\text{ are positive and radially }$
@@
We begin with the following straightforward observation: controlling the energy functional $\mathcal{F}$ from below is equivalent to bounding the mixed term $\int_{\Omega}uv$ from above.
@@
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R>0$ with some $n\geq 4$. Then
$\mathcal{F}(u,v)\leq\int_{\Omega}uv+\frac{\omega_{n}R^{n}}{e}$ (27)
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$, where $\omega_{n}$ is the surface area of the unit ball in $\mathbb{R}^{n}$.
@@
It follows from the basic inequality $s\ln s\geq-e^{-1}$ valid for all $s>0$. ∎
We incorporate decoupling techniques from Cao and Fuest [27], which assist us to convert the mixed term $\int_{\Omega}uv$ into an interior $L^{2}$–estimate for the Laplacian of the chemical concentration $v$. We will handle quadratic zeroth-order terms later.
@@
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R>0$ with some $n\geq 4$. There exists a constant $C_{2}>0$ such that
$\int_{\Omega}uv\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}+C_{2}(m+M)B\rho^{2-\kappa},$
@@
In view of (23), we test $u=-\Delta w+w$ by $v\mathds{1}_{B_{\rho}}$, and compute upon integration by parts,
$\begin{split}\int_{B_{\rho}}uv&=-\int_{B_{\rho}}\Delta wv+\int_{B_{\rho}}wv=-\int_{\partial B_{\rho}}v\partial_{\nu}w\;\mathrm{d}S+\int_{B_{\rho}}\nabla w\cdot\nabla v+\int_{B_{\rho}}wv\\ &=\int_{B_{\rho}}(-\Delta v+v)w+\int_{\partial B_{\rho}}(w\partial_{\nu}v-v\partial_{\nu}w)\;\mathrm{d}S=:J_{1}+J_{2}.\end{split}$ (28)
Using $w=-\Delta v+v-f$ specified in (21), we estimate by Young inequality
$\begin{split}J_{1}&=\int_{B_{\rho}}(-\Delta v+v)^{2}+\int_{B_{\rho}}(\Delta v-v)f\leq\frac{3}{2}\|\Delta v-v\|_{L^{2}(B_{\rho})}^{2}+\frac{1}{2}\|f\|_{L^{2}(B_{\rho})}^{2}\\ &\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}.\end{split}$ (29)
Recalling pointwise information from (26), we estimate
$J_{2}\leq MB\int_{\partial B_{\rho}}(|x|^{2-n}|x|^{1-\kappa}+|x|^{2-\kappa}|x|^{1-n})\;\mathrm{d}S=2\omega_{n}MB\rho^{2-\kappa},$ (30)
and also estimate the integral in the annular region
$\int_{\Omega\backslash B_{\rho}}uv\leq\sup_{r\in\left(\rho,R\right)}v(r)\cdot\int_{\Omega\setminus B_{\rho}}u\leq B\sup_{r\in\left(\rho,R\right)}r^{2-\kappa}\cdot\int_{\Omega}u=Bm\rho^{2-\kappa}.$ (31)
Summing (28)–(31) yields the desired estimate
$\int_{\Omega}uv\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}+(2\omega_{n}MB+Bm)\rho^{2-\kappa},$
for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ and $\rho\in(0,R)$, which completes the proof. ∎
To leverage the structural information (22), we employ the Pohožaev-type test function $|x|^{n-2}x\cdot\nabla v$ over the inner region $B_{\rho}$, initially introduced in the seminal work of Winkler [38, 10, 6]. This strategy enables us to establish an upper bound for the interior estimate $\|\Delta v\|_{L^{2}(B_{\rho})}$ through a combination of dissipation rate quantities and lower-order signal components. Specifically, this procedure generates a small factor $\rho^{2}$ in the critical term $\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2}$. Through careful calibration (see Lemma 4.5), we shall demonstrate that $\rho$ can be adaptively selected depending on $\|\nabla f\|_{L^{2}(\Omega)}$ to effectively convert this quadratic term into a sublinear power of the dissipation rate quantity $\|\nabla f\|_{L^{2}(\Omega)}^{2}$. Such test function methodologies have become standard tools in modern elliptic analysis, as evidenced by recent developments in [31].
@@
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R>0$ with some $n\geq 5$. Then there exists a constant $C_{0}>0$ such that
$\begin{split}&\frac{1}{8}\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+\frac{3}{4}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}\leq 12\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2}+C_{0}\|f\|_{L^{2}(B_{\rho})}^{2}\\ &+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}+\|v\|_{L^{2}(B_{\rho})}^{2}+C_{0}B^{2}\rho^{n+2-2\kappa}+C_{0}BM\rho^{2-\kappa}+2m\end{split}$ (32)
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ and $\rho\in(0,R)$.
@@
$-\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r+\int_{0}^{\eta}wv_{r}r^{\alpha}\;\mathrm{d}r$ $=\int_{0}^{\eta}(w_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r.$ (33)
@@
$\begin{split}\int_{0}^{\eta}(w_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r&=w_{r}(\eta)v_{r}(\eta)\eta^{\alpha}-\int_{0}^{\eta}w_{r}r^{n-1}(v_{r}r^{n-1})_{r}\;\mathrm{d}r\\ &=w_{r}(\eta)v_{r}(\eta)\eta^{\alpha}-\int_{0}^{\eta}w_{r}yr^{\alpha}\;\mathrm{d}r,\quad\eta\in(0,R).\end{split}$ (34)